OpenStudy (anonymous):

Hello dearies! Here is my problem... Solve for x: 9x^2-12x=-8 I got x=sqrt(8+12x)/3 Can you simplify this any more? The math brain is a little rusty haha

hartnn (hartnn):

do u know how to solve quadratic equations??

OpenStudy (australopithecus):

factor or use quadratic equation

OpenStudy (australopithecus):

never mind this cannot be factored nicely

OpenStudy (anonymous):

yes yes. it should be imaginary? and sorry i typed it wrong the 8 should be negative in my answer

OpenStudy (australopithecus):

Video on how to use the quadratic formula: https://www.youtube.com/watch?v=EeVqtpuMFOU

OpenStudy (anonymous):

write it in standard form 9x^2-12x+8=0 compare with ax^2+bx+c=0 here a=9 b=-12 c=8 now use Quadratic equation \[\Large x=\frac {-b \pm \sqrt{(b^2-4(a)(c))}}{2(2)}\]

OpenStudy (australopithecus):

yes there are no real solutions to this problem

OpenStudy (anonymous):

it is 2(a) in the denominator

OpenStudy (anonymous):

you will get an imaginary number here

OpenStudy (anonymous):

oh i am silly. i had convinced myself that you can't use quadratic formula to find imaginary answers...thank you loves!

OpenStudy (anonymous):



OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yup just reduce